OpenStudy (anonymous):

Marks: 5 [10.07] What is the exact value of 8 times z squared all over the square root of z to the seventh power.? Choose one answer. a. 8 the square root of z all over z squared. b. 8 c. 8 times z over the square root of z. d. 8 the square root of z all over z.

OpenStudy (allank):

Rewriting in symbol form should make it simpler (8*z^2)/sqroot(z)^7

OpenStudy (anonymous):

thanx

OpenStudy (campbell_st):

simplify the denominator \[\frac{8z^2}{\sqrt{z} \times \sqrt{z^6}} = \frac{8z^2}{z^3\sqrt{z}}\] which simplifies to \[\frac{8}{z \sqrt{z}}\] looking at your choice you may need to rationalise the denominator

OpenStudy (anonymous):

thanks for the help guys